# Physics-informed DeepCT: Sinogram Wavelet Decomposition Meets Masked Diffusion

### v1

@@ 3.2.1 Random Mask Strategy in Sinogram Domain
In the prior learning phase based on the projection domain, we propose a training strategy that involves embedding random masks into sinograms. By applying random masks to the sinograms, this method introduces random perturbations within the limited data sample space, effectively increasing the diversity of the feature space and enhancing the model focus on a wider range of data distributions, thereby improving its generalization ability as well as the reliability and accuracy of the reconstructed results.
According to the physical characteristics of projection data acquisition, we have specially designed a random mask strategy to address issues such as missing scanning angles, detector ray damage, and the random loss of some projection data. By applying randomly generated masks to the projection data, we introduce uncertainty to enhance model robustness. Specifically, given full-view projection data $y$ and the random mask operator $\hat{m}$, the masked full-view projection data can be expressed as:
@@
where $\tilde{Y}_{s}$ indicates the projected data with the mask.
The mask strategy as follows: Given arbitrary mask type, the seed set $M=\{M_{1},M_{2},M_{3}\}$. The type symbol $S$ is generated randomly from $\{1,2,\cdots,n\}$. Randomly generated mask type symbol is marked as $M_{s}$. The set of random mask locations under $M_{s}$ is marked as $\hat{S}$. The randomly mask generated as follows:
$\hat{m}_{i,j}=\begin{cases}0,&(i,j)\in\hat{S},\\ 1,&(i,j)\notin\hat{S},\end{cases}$ (9)
where $s\sim Uniform\{1,2,\cdots,n\}$ , $\hat{m}$ is the mask generated by random rules.
3.2.2 Random High-frequency in Sinogram Wavelet Domain
Prior learning in the projection domain focuses more on global information and and tends to ignore the details. To improve the detail quality of reconstructed images, we propose a sinogram wavelet-based random high-frequency diffusion strategy. By performing wavelet decomposition of the sinogram and randomly selecting high-frequency components as network inputs for training, this approach enables the model to focus on specific frequency details. Additionally, the introduction of random high-frequency components enhances feature representation in the data space. This improvement bolsters the model’s adaptability to a variety of frequency features and significantly aids in the recovery of fine textures and other detailed characteristics in the reconstructed images.
Specifically, decomposing the sinogram into four sub-images using Discrete Wavelet Transform (DWT): the low-frequency approximation component (LL) and the high-frequency detail components (LH, HL, HH). This process can be expressed as:
$H:y\rightarrow\{h_{LL}(y),h_{LH}(y),h_{HL}(y),h_{HH}(y)\},$ (10)
where $H$ represents the wavelet transform, $h_{LL}$ represents the low-frequency component after the transformation, $h_{LH}$, $h_{HL}$ and $h_{HH}$ correspond to the high-frequency detail components in the vertical, horizontal and diagonal directions respectively. Let the three high-frequency components be denoted as $Y_{h}=\{h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$. By randomly selecting a high-frequency sub-image $\tilde{Y}_{h}$ from the wavelet domain for training, the ability of the network to learn texture details.
3.2.3 Forward Diffusion Process Based on SDEs
The two-stage training is carried out through a score generation model based on stochastic differential equations (SDEs) to obtain the Sinogram Random Mask model (SRM) and Sinogram wavelet random High-frequency Deocomposition model (SHD), respectively. In the training stage, the model mainly includes two processes: the forward process and the backward process. As shown in Fig. 2(a) and Fig. 2(b), the forward process of the projection domain and the wavelet domain during the training phase is illustrated in detail. In this process, the forward stochastic differential equation (SDE) progressively transforms a complex data distribution into a tractable Gaussian distribution by gradually adding noise.
Given a diffusion process $\{x(t)\}^{T}_{t=0}$, represented by a continuous time variable $t\in[0,T]$, such that $x(t)\sim p_{t}$, where the sample image data is independently and identically distributed. The forward SDE process is expressed as follows[30]:
$dx=f(x,t)dt+g(t)dw,$ (11)
where $f(x,t):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is a vector-valued function and $\mathbb{R}\rightarrow\mathbb{R}$ is a scalar function about $x(t)$, which are called drift coefficient and diffusion coefficient respectively. $N$ indicates the dimension of the image data in the projection domain. $w\in\mathbb{R}^{N}$ induces the standard Brownian motion.
Variance explosion SDE (VE-SDE) is obtained by applying $f(x,t)=0$, $g(t)=\sqrt{d[\sigma^{2}(t)]/dt}$ in order to improve the capability. In this case, the two-stage VE-SDE is represented as follows:
$d\tilde{Y}_{s}=\sqrt{d[\sigma^{2}(t)]/dt}dw,$ (12)
$d\tilde{Y}_{h}=\sqrt{d[\sigma^{2}(t)]/dt}dw,$ (13)
where $\sigma(t)>0$ represents a monotonically increasing function and signifies the time-varying escalating scale function for noise.
@@ 3.2.3 Forward Diffusion Process Based on SDEs
where $\lambda_{t}$ is a positive function, $\log p_{0t}(\tilde{Y}_{s}(t)|\tilde{Y}_{s}(0))$ is the gaussian perturbation kernel centered at $\tilde{Y}_{s}(0)$. Once the network satisfies $\tilde{Y}_{s}(t)\simeq\nabla_{\tilde{Y}_{s}(t)}\log p_{0t}$ and $\nabla_{\tilde{Y}_{s}(t)}\log p_{0t}$ will be known for all $t$ by solving $\nabla_{\tilde{Y}_{s}(t)}\log p_{0t}$.
The prior learning stage for random wavelet high-frequency components involves training the high-frequency components through the neural network $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h}(t),t)$. Gaussian noise is gradually introduced into randomly selected high-frequency wavelet components to optimize the network parameters. The optimization objective function of this process is as follows:
@@
Alg. 1 provides a detailed description of SRM and SHD in the training process.
Algorithm 1
SRM Training process
1:  Load projected data $y$ from the training set;
2:  Randomly generated a discrete random mask $\hat{m}$;
3:  Generate a masked sinogram: $\tilde{Y}_{s}=y\odot\hat{m}$;
4:  Trained with SRM;
5:  Output $\textbf{s}_{\theta_{s}}(\tilde{Y}_{s},t)$.
SHD Training process
1:  Load projected data $y$ from the training set;
2:  Wavelet transform:
    $H:y\rightarrow\{h_{LL}(y),h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$;
3:  Random selection of high-frequency components:
     $\tilde{Y}_{h}\leftarrow\{h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$;
4:  Trained with SHD;
5:  Output $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h},t)$.
To sum up, $\textbf{s}_{\theta_{s}}(\tilde{Y}_{s}(t),t)$ and $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h}(t),t)$ can accurately learn the data distribution of real images in the projection domain and the high-frequency data distribution of sinogram wavelet. This approach not only enhances the model ability to recognize the global structure of the image but also improves its sensitivity to details such as texture. With this dual-model strategy, the model is further encouraged to learn and approximate the underlying probability distribution of the data.
3.3 Cascade Reconstruction of Sinogram and Wavelet
The image generation process in iterative reconstruction is essentially the inverse problem of SDEs, specifically a reverse diffusion process. By utilizing prior knowledge from the SRM and SHD, we ensure that the reconstruction stage generates comprehensive global information and finer details. Combined with data consistency constraints, SDEs employ a predictor-corrector (PC) sampler to efficiently generate high-quality images. This process can be represented by the following reverse-time SDE:
@@ 3.3 Cascade Reconstruction of Sinogram and Wavelet
In the joint iterative reconstruction stage, by introducing regularized prior information, the model is guided to gradually integrate the global consistency in the projection domain with the detailed features in the high-frequency components of the wavelet. The overall optimization objectives are as follows:
@@
where $\tilde{H}[\cdot]$ represents the full-frequency information, $\tilde{H}_{l}[\cdot]$ represents the low-frequency component and $\tilde{H}_{h}[\cdot]$ represents the high-frequency component.
We optimize the two-stage process through successive iterations of SRM and SHD. Specifically, SRM is optimized using a sinogram-based score model, while SHD is optimized by enhancing the details of the high-frequency components in the wavelet domain. This approach facilitates the transition from global consistency to fine details. The process can be formalized into the following two steps:
$\hat{y}_{s}^{t-1}=\underset{\hat{y}_{s}}{\arg\min}\ \|P(\wedge)y-\hat{y}_{s}\|^{2}_{2}+\nu_{s}R_{s}(\hat{y}_{s}^{t}),$ (20)
$\hat{y}_{h,i}^{t-1}=\underset{\hat{y}_{h,i}}{\arg\min}\ \|\tilde{H}_{h}(P(\wedge)y)-\hat{y}_{h,i}\|^{2}_{2}+\nu_{h}R_{h}(\tilde{H}_{h}[\hat{y}_{s}^{t-1}]).\vspace{-0.5 em}$ (21)
3.3.1 Sinogram Generation
The sinogram generation stage, as the first subproblem, is designed to enhance global information to ensure data consistency and integrity. This stage can be broken down into the following two subproblems:
@@ 3.3.1 Sinogram Generation
Eq. (22) can be considered a data consistency term for the sinogram, which can be simplified to the following equation by ensuring the alignment between the generated data and the original data:
$\hat{y}_{s}^{t-\frac{1}{2}}=(1-P(\wedge))\hat{y}_{s}^{t-\frac{1}{2}}+P(\wedge)y.\vspace{-0.5 em}$ (24)
The predictor serves as a variance explosion (VE) SDE solver to generate an estimate for the next reconstruction iteration, which is solved numerically through backdiffusion. Using the trained model $\textbf{s}_{\theta_{s}}$ for the reverse SDE generates samples, and the discretization of the predictor is formulated as follows:
@@
3.3.2 High-frequency generation in wavelet domain
Upon completion of the reconstruction in the projection domain during the first stage, the sinogram is subsequently transformed into the wavelet domain. In this domain, three high-frequency components are extracted, whereas the low-frequency component is preserved. The process can be described as follows:
@@ 3.3.2 High-frequency generation in wavelet domain
Building on the approach used for the subproblem in the projection domain during the first stage, the high-frequency components extracted from the wavelet domain are employed to optimize the score-based generation model. This optimization refines intricate details, such as complex textures. The updated objective function is expressed as follows:
$\hat{y}_{h,i}^{t-\frac{1}{2}}=\underset{\hat{y}_{h}}{\arg\min}\ \|\hat{y}_{h,i}-\tilde{H}_{h}[P(\wedge)y]\|^{2}_{2}+\mu_{h}\|\hat{y}_{s}^{t-1}-\hat{y}_{s}\|^{2}_{2},$ (28)
$\hat{y}_{h,i}^{t-1}=\underset{\hat{y}_{h}}{\arg\min}\ \|\hat{y}_{h,i}-\hat{y}_{h,i}^{t-\frac{1}{2}}\|^{2}_{2}+\nu_{h,i}R_{h}(\hat{y}_{h,i}),$ (29)
where $\hat{y}_{h,i}$ represents the $i$-th high-frequency component and $i=0,1,2$.
According to the data consistency of the high-frequency components in the wavelet domain, to ensure the reliability of the data:
$\hat{y}_{h,i}^{t-\frac{1}{2}}=\tilde{H}_{h}[(1-P(\wedge))\hat{y}_{s}^{t-\frac{1}{2}}+P(\wedge)y].$ (30)
At this stage, the same “predictor” and “corrector” steps of the VE-SDE diffusion model are applied to the high-frequency components of each channel. Specifically, the “predictor” step uses the trained $\textbf{s}_{\theta_{h}}$ network for reconstruction, as follows:
@@
3.3.3 Domain transform stage
Ultimately, the process integrates the retained low-frequency information with the optimized high-frequency components. This combined data is then transformed back to the projection domain via wavelet inversion, as detailed below:
@@ 3.3.3 Domain transform stage
In summary, the iterative reconstruction phase of SWARM is illustrated in Fig. 2(c). The models are trained by embedding random masks in the sinogram and employing random high-frequency strategies in the sinogram wavelet domain to capture the prior distributions of global information and detailed features. During reconstruction, the numerical SDE solver and Langevin dynamics are iteratively updated to obtain high-quality full-view projection data. The detailed process of the reconstruction stage is shown in Alg. 2.
Algorithm 2
SWARM Iterative Reconstruction
@@
4:   Update $\hat{Y}_{s}^{t-\frac{1}{2}}$ by data consistency with Eq. (24);
@@
6:   Update $\hat{Y}_{s}^{t-1}$ by data consistency with Eq. (24);
@@
9:     Update $\hat{Y}_{h,i}^{t-\frac{1}{2}}$ by data consistency with Eq. (30);
@@
11:     Update $\hat{Y}_{h,i}^{t-1}$ by data consistency with Eq. (30);
@@
4 EXPERIMENT
4.1 Data Specification
4.1.1 AAPM Challenge Data

### v2

@@ III-B Training Process in Sinogram Wavelet Domain
1) Prior Learning for Random Mask Strategy in Sinogram Domain: In the prior learning stage based on the projection domain, we propose a virtual random mask training strategy. Combining the physical characteristics of projection data acquisition and the randomness of embedding masks in the projection domain. The Sinogram Random Mask model (SRM) is constructed through the SDEs to enhance the robustness of the model against distribution offsets. As shown in Fig. 2(a), by applying random masks to the sinograms, SRM introduces random perturbations within the limited data sample space. It effectively augments the diversity of the feature space and sharpens the model focus on a broader spectrum of data distributions. This improves the generalization ability of the model as well as the reliability and accuracy of the reconstruction results.
To tackle challenges such as limited scanning angles, we have devised a specialized random masking strategy. By defining a probability parameter $p\in[0,1]$, this strategy is applied to simulate sparse-view sampling scenarios through stochastic masking of projection data. Specifically, for each training sample: a sparse-view mask is applied with probability $p$. In this case, one masking pattern is randomly selected from a predefined set of sparse-view configurations. With probability $1-p$, no masking is applied, and the full projection data is retained. Given a full-view projection data $y$ and the random mask operator $\hat{m}$, the masked full-view projection data can be expressed as:
@@
where $\tilde{Y}_{s}$ indicates the projected data with the mask. This probabilistic masking mechanism introduces controlled randomness, enhancing the model’s robustness to varying degrees of data sparsity and promoting generalization across different undersampling scenarios.
The forward stochastic differential equation (SDE) progressively transforms a complex data distribution into a tractable gaussian distribution by gradually adding noise. Given a diffusion process $\{y(t)\}^{T}_{t=0}$, represented by a continuous time variable $t\in[0,T]$, such that $y(t)\sim p_{t}$, where the sample image data is independently and identically distributed. The general forward process of SDE is expressed as follows[31]:
$dx=f(y,t)dt+g(t)dw,$ (9)
where $f(y,t):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is a vector-valued function and $\mathbb{R}\rightarrow\mathbb{R}$ is a scalar function about $y(t)$, which are called drift coefficient and diffusion coefficient respectively. $N$ indicates the dimension of the image data in the projection domain. $w\in\mathbb{R}^{N}$ induces the standard Brownian motion. Variance explosion SDE (VE-SDE) is obtained by applying $f(y,t)=0$, $g(t)=\sqrt{d[\sigma^{2}(t)]/dt}$ in order to improve the capability.
In this case, the VE-SDE is represented as follows:
$d\tilde{Y}_{s}=\sqrt{d[\sigma_{s}^{2}(t)]/dt}dw_{s},$ (10)
where $\sigma_{s}(t)>0$ represents a monotonically increasing function and signifies the time-varying escalating scale function for noise.
@@
where $\lambda_{t}$ is a positive function, $\log p_{0t}(\tilde{Y}_{s}(t)|\tilde{Y}_{s}(0))$ is the gaussian perturbation kernel centered at $\tilde{Y}_{s}(0)$. Once the network satisfies $\textbf{s}_{\theta_{s}}(\tilde{Y}_{s}(t),t)$ and $\nabla_{\tilde{Y}_{s}(t)}\log p_{0t}$ will be known for all $t$ by solving $\nabla_{\tilde{Y}_{s}(t)}\log p_{0t}$.
2) Prior Learning for Random High-frequency in Sinogram Wavelet Domain: In order to improve the detail quality of the reconstructed image, a random high-frequency training strategy based on sinogram wavelets is proposed to obtain the Sinogram wavelet random High-frequency Decomposition model (SHD). In this approach, the projection data is decomposed into frequency bands using orthogonal wavelet basis functions. A dynamic random sampling mechanism is then formulated for high-frequency components, where learnable weights are employed to quantify the salience of each sub-band. By leveraging learnable weights, the significance of each sub-band is assessed, enabling the generation of training data with anisotropic perturbations in the frequency domain. Furthermore, this method clearly distinguishes the high-frequency local details in the projection domain. This forces the network to adaptively focus on the information of key frequency bands.
Specifically, decomposing the sinogram into four subbands using wavelet transform (DWT): the low-frequency approximation component (LL) and the high-frequency detail components (LH, HL, HH). This process can be expressed as:
$H:y\rightarrow\{h_{LL}(y),h_{LH}(y),h_{HL}(y),h_{HH}(y)\},$ (12)
where $H$ represents the wavelet transform, $h_{LL}$ represents the low-frequency component after the transformation, $h_{LH}$, $h_{HL}$ and $h_{HH}$ correspond to the high-frequency detail components in the vertical, horizontal and diagonal directions respectively. Let the three high-frequency components be denoted as $Y_{h}=\{h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$. Training with the dynamic features $\tilde{Y}_{h}$ obtained through random sampling of the high-frequency sub-bands in the wavelet domain effectively enhances the network’s ability to represent high-frequency details. The VE-SDE is represented as follows:
$d\tilde{Y}_{h}=\sqrt{d[\sigma_{h}^{2}(t)]/dt}dw_{h},$ (13)
where $\sigma_{h}(t)>0$ represents a monotonically increasing function and signifies the time-varying escalating scale function for noise. The prior learning stage for random wavelet high-frequency components involves training the high-frequency components through the neural network $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h}(t),t)$. Gaussian noise is gradually introduced into the randomly selected high-frequency sub-bands in the wavelet domain to optimize the network parameters. The optimization objective function of this process is:
@@
With sufficient data and model capacity, score matching ensures that $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h}(t),t)\approx\nabla_{\tilde{Y}_{s}(t)}\log p_{0t}$ for almost all $\tilde{Y}_{h}$ and $t$.
To sum up, $\textbf{s}_{\theta_{s}}(\tilde{Y}_{s}(t),t)$ and $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h}(t),t)$ can accurately learn the data distribution of real images in the projection domain and the high-frequency data distribution of sinogram wavelet. This method not only enhances the model’s ability to identify the overall structure of the image but also improves its sensitivity to details such as textures. By adopting this dual-model strategy, the model is further encouraged to learn and approximate the underlying probability distribution of the data.
III-C Cascade Reconstruction of Sinogram and Wavelet
In iterative reconstruction, the image generation process can essentially be formulated as the inverse problem of SDEs. By incorporating prior knowledge from SRM and SHD, the reconstruction phase captures not only comprehensive global structures but also preserves fine-grained local details. Leveraging these priors along with data consistency constraints, the SDE framework employs a Predictor-Corrector (PC) sampler to enable stable and efficient reverse-time evolution from noise to data distribution, thereby producing high-quality images. Reverse-time SDE can be characterized as follows:
@@ III-C Cascade Reconstruction of Sinogram and Wavelet
During the joint iterative reconstruction stage, the introduction of regularized prior information effectively guides the model to progressively integrate the global consistency features from the projection domain with the fine-grained details embedded in the high-frequency components of the wavelet domain, thereby enabling the collaborative restoration of structural and textural information. The overall optimization objective can be formulated as follows:
@@
where $\tilde{H}[\cdot]$ represents the full-frequency information, $\tilde{H}_{l}[\cdot]$ represents the low-frequency component and $\tilde{H}_{h}[\cdot]$ represents the high-frequency component. Through the iterative optimization mechanism of SRM and SHD, a smooth transition from global consistency to detail precision is effectively achieved.
1) Sinogram Generation: The sinogram generation stage enhances the global structural information in the image, thereby ensuring the consistency and integrity of the data during the processing. In this stage, it can be further refined into two interrelated subtasks, which respectively focus on the modeling of global information and the expression and optimization of structural features, so as to improve the expressive ability of the overall data and the reconstruction accuracy. This stage can be decomposed into the following two sub-problems:
@@
On the basis of ensuring the consistency between the generated data and the original observed data, the data consistency constraint term enhances the stability of the model optimization process.
$\hat{y}_{s}=(1-P(\wedge))\hat{y}_{s}+P(\wedge)y.\vspace{-0.5 em}$ (21)
The predictor serving as a solver for the Variance Exploding type of Stochastic Differential Equation, generates estimates for updates in each reconstruction iteration and numerically solves this SDE through the reverse diffusion process. In this process, the pre-trained model $\textbf{s}_{\theta_{s}}$ is utilized to sample the reverse SDE, thereby achieving the gradual generation of samples. The discretization process can be expressed as follows:
@@
2) High-frequency generation in wavelet domain: After the preliminary reconstruction is completed in the projection domain during the first stage, the obtained sinogram is mapped to the wavelet domain. In this domain, the image is decomposed at multiple scales, from which the high-frequency components in three directions are extracted, and the low-frequency components are retained simultaneously, so as to fully capture the detailed structures and the overall contours in the image. The process can be described as follows:
@@
The data consistency of the high-frequency components in the wavelet domain to ensure the reliability of the data:
$\hat{y}_{h,i}=\tilde{H}_{h}[(1-P(\wedge))\hat{y}_{s}+P(\wedge)y].$ (25)
The “predictor-corrector” framework in the VE-SDE is also applied to the reconstruction process of each high-frequency channel. Specifically, in the “predictor” step, the network $\textbf{s}_{\theta_{h}}$ is utilized to reconstruct each high-frequency component, and the process can be described as follows:
@@
3) Domain transform stage: Finally, this process integrates the previously retained low-frequency information with the optimized high-frequency components. By fusing information across different frequency levels, a more comprehensive image representation is achieved. Subsequently, the combined data is mapped back to the projection domain through the inverse wavelet transform, as detailed below:
@@
In summary, the iterative reconstruction stage of SWARM is shown in Fig. 2(c). By embedding a random mask and introducing a random high-frequency strategy in the sinogram, the prior distributions of global structural and detailed features are extracted, effectively guiding the training process of the model. In the actual reconstruction, through the iterative updates of the numerical SDE solver and the Langevin dynamics, high-quality full-view projection data is gradually obtained. The specific reconstruction process is shown in Fig. 2.
Alg. 1 provides a detailed description of SRM and SHD in the training process. The detailed process of the reconstruction stage is shown in Alg. 2.
Algorithm 1 Training process.
   Dataset: Load image data, generate projection data via FP forward projection.
   SRM Training process:
   1:  Sample a projected data $y$;
   2:  Randomly generated a discrete random mask $\hat{m}$;
   3:  Generate a masked sinogram: $\tilde{Y}_{s}=y\odot\hat{m}$;
   4:  Trained with SRM;
   5:  return $\textbf{s}_{\theta_{s}}(\tilde{Y}_{s},t)$.
   SHD Training process:
   1:  Sample a projected data $y$;
   2:  Wavelet transform:
   3:  $H:y\rightarrow\{h_{LL}(y),h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$;
   4:  Random selection of high-frequency components:
   5:  $\tilde{Y}_{h}\leftarrow\{h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$;
   6:  Trained with SHD;
   7:  return $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h},t)$.
Algorithm 2 Iterative Reconstruction process.
@@
   4:     Update $\hat{Y}_{s}^{t-\frac{1}{2}}$ with Eq. (21);
@@
   6:     Update $\hat{Y}_{s}^{t-1}$ with Eq. (21);
@@
   9:        Update $\hat{Y}_{h,i}^{t-\frac{1}{2}}$ with Eq. (25);
@@
   11:       Update $\hat{Y}_{h,i}^{t-1}$ with Eq. (25);
@@
IV Experiment
IV-A Data Specification